John is 28 years older than his son. In two years he will be five times older than his son. How old are both of them now?
Answer by ewatrrr(24785)

Now
Let x represent son's age
John (x+28)
In two years...
(x+28) + 2 = 5(x+2)
x + 30 = 5x + 10
20 = 4x
5 = x , son's age NOW. John is 33.
And..in two years...
35 = 5*7 Checks!